# The Multiband Imaging Survey for High-Alpha PlanetS (MISHAPS) I: Preliminary Constraints on the Occurrence Rate of Hot Jupiters in 47 Tucanae

### v1

@@ Abstract
facilities: Blanco (DECam)software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), CFITSIO (Pence, 1999), DoPHOT (Mateo & Schechter, 1989; Schechter et al., 1993), FITSH (Pál, 2012), ISIS (Alard & Lupton, 1998; Alard, 2000), JKTEBOP (Southworth, 2013), JKTLD (Southworth, 2008), matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020), Source Extractor (Bertin & Arnouts, 1996), VARTOOLS (Hartman & Bakos, 2016), WCSTOOLS (Mink, 1998, 2019)
1 Introduction
Though giant planet formation rates have been found to correlate with [Fe/H] (e.g., Gonzalez, 1997; Santos et al., 2004; Fischer & Valenti, 2005; Johnson et al., 2010), the specific elements responsible for the correlation have largely been unresolved. Volatiles and silicates are expected to be crucial ingredients for giant planet formation, because the silicate dust grains in protoplanetary disks provide surfaces for the more abundant volatiles to stick to, forming planetesimals (Pontoppidan et al., 2014; Öberg & Bergin, 2021; Nielsen et al., 2023). Beyond the ice line, these volatiles likely make up the majority of the protoplanetary disk’s solid mass. These compounds largely consist of $\alpha$-elements — elements formed in $\alpha$-capture reactions — such as oxygen, silicon, and magnesium, suggesting that the $\alpha$-abundance, [$\alpha$/H] = [Fe/H] + [$\alpha$/Fe], may influence planet occurrence rates more than [Fe/H]. However, our current sample of giant planet hosts prevents us from disentangling a relationship. Many come from Kepler and TESS, whose targets primarily lie within the high-[Fe/H], low-[$\alpha$/Fe] thin disk (Bensby et al., 2003, 2005; Reddy et al., 2006; Guo et al., 2017; Weinberg et al., 2019; Griffith et al., 2021). Past observational limits for ground-based radial velocity and transit searches have similarly limited our sample to primarily disk stars. Some analyses of low-[Fe/H], high-[$\alpha$/Fe] subsamples of stars in the Galactic thick disk such as Adibekyan et al. (2012a, b) have found higher numbers of giant planets than are predicted by a [Fe/H] dependence, which could potentially be explained by the relative $\alpha$-enhancement. However, the bias of our giant planet sample towards disk stars prevents us from determining a conclusive explanation for this surplus: we lack observations in the full [$\alpha$/Fe]-[Fe/H] parameter space, particularly in high-[Fe/H], high-[$\alpha$/Fe] populations (e.g., Wilson et al., 2022).
Additionally, globular clusters allow us to search for planets in dense stellar environments — a region of parameter space that is largely unexplored. This parameter space is complicated by some theoretical uncertainty: we don’t yet know if, e.g., the gravitational interactions produced by these dense fields would inhibit or enhance hot Jupiter formation. Hamers & Tremaine (2017) posit that only the densest core regions of the clusters would eject all planets, and Winter et al. (2020) suggest that the gravitational perturbations in dense fields may initiate planet-planet scattering that actually enhances hot Jupiter formation. Other studies suggest that conditions in globular clusters beyond gravitational effects may adversely affect the likelihood of planet formation. Thompson (2013) theorizes that because dense regions such as globular clusters tend to form in hot starbursts, the high temperatures may evaporate the protoplanetary disks more quickly than giant planets can form. Further, Lee & Hopkins (2020) find that the ambient radiation in star forming regions is extremely high, which could have yet-uncharacterized effects on planet formation, such as potentially eliminating ice lines in the protoplanetary disks entirely. Thus, tighter observational constraints on planet occurrence rates in extreme environments are needed to constrain these seemingly contradictory theories.
To date, planet searches of globular clusters have resulted in one successful detection in M4: a roughly 1.18 $R_{\text{Jup}}$ planet orbiting the millisecond pulsar PSR B1620-26 (Richer et al., 2003). Its location towards the core of M4 implies that planet formation in clusters may still be possible, despite their complex stellar dynamics (Fregeau et al., 2006). Other studies suggest it may indicate that — though globular clusters may have too low a metallicity for typical giant planet formation — they may be suitable for forming planets from supernova fall-back (e.g., Beer et al., 2004; Sigurdsson et al., 2008). However, no searches for globular cluster planets orbiting normal main sequence or giant stars have successfully discovered a planet, including a further search of M4 using K2 data (Wallace et al., 2020), which was limited to only $\sim$4,000 distinguishable stars due to Kepler’s 4”/px resolution.
These unsuccessful searches include surveys of 47 Tucanae and $\omega$ Centauri, some of the most massive and closest clusters to Earth. 47 Tuc in particular has been promoted as a potential planet-hosting cluster, as analyses of pulsar planet dynamics such as Sigurdsson (1992) place its density in a dynamical sweet spot where a giant planet could form from supernova ejecta remaining around a pulsar before the material is disrupted by stellar interactions. Three different studies, inluding two surveys, have therefore assessed 47 Tuc’s planet-hosting abilities (Gilliland et al., 2000; Weldrake et al., 2005; Masuda & Winn, 2017). $\omega$ Cen was searched by Weldrake et al. (2008), which analyzed $\sim$31,000 main sequence stars in the cluster, finding no planets. An additional survey by Nascimbeni et al. (2012) analyzed over 5,000 lightcurves from HST observations of another globular cluster, NGC 6397, focusing particularly on a sample of 2,215 M-dwarfs, and again found no planets.
@@ 1 Introduction
Surveys of 47 Tuc began with an HST search by Gilliland et al. (2000) (hereafter G00). Using the accepted hot Jupiter occurrence rate in the solar neighborhood for the time — 0.8-1.0% — they expected to find $\sim$17 planets in their observations. However, despite taking 8.3 days of near-continuous data for $\sim$34,000 stars, they observed no statistically significant transit events. In a later re-analysis using a simulated stellar sample based on giant planet hosting stars in the Kepler sample, Masuda & Winn (2017) (hereafter MW17) find that, if the occurrence rate of hot Jupiters in 47 Tuc was to match that of Kepler field stars, accounting for the host-mass dependence of giant planet occurrence rate (Johnson et al., 2010), G00 could only have expected to find $\sim$2 planets in their sample of 34,000 stars, with a $\sim$15chance of finding no planets. Thus, the null result of the G00 survey is less statistically significant than initially thought.
@@
Weldrake et al. (2005) (hereafter W05) followed-up the G00 survey with a 33-night ground-based search using the Australian National University 40 inch Telescope. To complement the G00 search, which focused more on the core of the cluster, this survey focused on $\sim$22,000 stars in the outskirts of the cluster. Using Monte Carlo simulations, they expected to find $\sim$7 planets in their data if the occurrence rate was 0.8%, and still found none.
@@ 2.1 Survey Design
Motivated by MW17, our goals are to place stronger constraints on the occurrence rate than G00’s observations allow, and to conclusively rule out an occurrence rate equal to that in the Kepler field (average [M/H]=-0.045$\pm$0.009, e.g., Guo et al. (2017)).
MISHAPS uses the Dark Energy Camera (DECam) on the 4-m Víctor M. Blanco telescope at Cerro-Tololo Interamerican Observatory to perform a high-cadence, wide-field survey of 47 Tuc and other $\alpha$-enhanced fields to search for transiting planets. DECam’s wide field of view and the Blanco’s 4-m diameter allow us to pursue an alternative strategy to most transiting planet searches that have relied in part on covering multiple transits and folding to increase the signal-to-noise ratio (S/N) above a detection threshold. Instead, the Blanco’s aperture enables a clear detection of a Jupiter-radius planet’s transit over $\sim$5 magnitudes of the 47 Tuc main sequence. We therefore planned a small survey ($\sim$7 nights) to search for single full or partial transits on a larger sample of stars than either G00 or W05. Without the necessity of multiple transit detections, the detection probability for short-period planets is therefore $\sim$10’s of percent for even 7 nights of data. With the expectation that the number of planet candidates we would find would be small, and that the quality of data would allow many EB false positives to be rejected from survey data alone, we anticipated that the number of targets that required follow-up would be small. For those that did, an estimate of the orbital period from a single transit would limit the amount of observations needed to recover a second transit and measure a period (e.g., Seager & Mallén-Ornelas, 2003; Yee & Gaudi, 2008; Villanueva et al., 2019).
We selected DECam for its size, photometric precision, near-infrared $z-$band sensitivity, and low overhead. DECam consists of 62 deep-depletion science CCDs, each of which has a size of 2048x4096 px for a total of 520 Mpx (Honscheid & DePoy, 2008), though only 60 are currently in use. The $\sim$3 deg2 FOV allows us to simultaneously monitor the full expanse of 47 Tuc. Its photometric precision allows us to resolve stars with magnitudes down to $r\approx 24$ and to potentially detect single transits on stars down to $r\approx 22$. DECam’s extremely high $z-$band sensitivity and low overheads are critical to enabling chromatic timeseries observations, which allow us to use differences in eclipse depths in $r$ and $z$ to rule out false positives caused by secondary eclipses of eclipsing binaries (EBs), diluted primary eclipses of hierarchical triple stars, and blended Small Magellanic Cloud (SMC) EBs. The low overheads also allow observations at a high enough cadence to resolve ingresses and egresses of transits, which are important for estimating accurate transit times and host stellar densities (e.g., Seager & Mallén-Ornelas, 2003), as well as increasing the accuracy with which one can estimate periods by adding well-sampled partial eclipses.
@@ 2.2 Observations
Preliminary observations of 47 Tuc were taken in short 2-4 hour windows on 17 nights in 2019 and 2021. Additionally, eight half-nights of observations of the cluster were conducted in 2022. Each night, we took a continuous timeseries of images with 100 s exposure times, alternating between the $r$ and $z$ filters with each exposure. This exposure sequence resulted in a $\sim 4$-minute cadence in each filter, or $\sim 2$ minutes combined. These sequences were interrupted roughly once per hour to check and correct the telescope pointing if it drifted by more than 10”. All images were first processed through the DECam Community Pipeline (Valdes et al., 2014), which applies calibration frames, corrects instrumental signatures, and applies a cosmic ray removal algorithm. We performed a preliminary determination of the overall quality of the nights by plotting the background, ellipticity, zeropoint magnitude, and seeing values from the image headers and flagging any nights which consistently had a zeropoint $<$30.0, the magnitude of a star resulting in only 1 photon s-1 in the exposure. These poor-quality nights were later confirmed via analysis of the nightly median absolute deviation (MAD) of the lightcurve’s photometric root-mean-square (RMS) after the images were processed and the photometry was extracted (§4.1). Table 2.2 shows the log of observing nights and their mean data quality statistics. In total, our observations produce 1,288 images in the $z-$band and 1,293 images in the $r-$band.
@@ 3 Reference Image Analysis
We produce reference images for each band for use in difference imaging, requiring us to choose an appropriate number of high-quality images from our full dataset and stack them (§3.1). We then extract our initial catalog of stars using PSF photometry, and find astrometric solutions and photometric calibrations for each target by crossmatching with Gaia DR3 (Babusiaux et al., 2023) and the NOIRLab Source Catalog Data Release 2 (NSC, Nidever et al., 2018, 2021)3.2). We also determine appropriate proper motion and color selections for cluster membership using the Gaia crossmatch results (§3.3). We then perform several characterization steps to determine which targets will be used to perform transit injections and for use in efficiency estimates (§3.4).
@@ 3.1 Reference Image Construction
For each of DECam’s CCD chips, we constructed a deep, high-quality reference image in the $r$ and $z$ filters by selecting and combining a small number of the best images from the data taken in 2021, which had the best conditions. To do this, the community pipeline-determined point spread function (PSF), full width at half maximum (FWHM), sky background, zeropoint (magnitude resulting in 1 photon s-1), PSF ellipticity, and airmass for all images were compiled from the FITS headers, and sorted by FWHM. Fifteen images were selected in order of increasing FWHM, and any with high sky background, PSF ellipticity, or smaller zeropoint magnitude were skipped. Each candidate image was visually inspected to ensure it was a good-quality image. The fifteen selected images were combined using the ISIS image subtraction program’s ref.csh task (Alard & Lupton, 1998; Alard, 2000), which aligns the images; determines a PSF-matching convolution kernel and photometric scaling to match the PSF, zeropoint, and sky background of the input images; and then combines them as a mean stack with 3-$\sigma$ clipping. We refer to the resultant images as the reference image for the relevant chip and filter.

### v2

@@ Abstract
The first generation of transiting planet searches in globular clusters yielded no detections, and in hindsight, only placed occurrence rate limits slightly higher than the measured occurrence rate in the higher-metallicity Galactic thick disk. To improve these limits, we present the first results of a new wide field search for transiting hot Jupiters in the globular cluster 47 Tucanae. We have observed 47 Tuc as part of the Multiband Imaging Survey for High-Alpha Planets (MISHAPS). Using 24 partial and full nights of observations taken with the Dark Energy Camera on the 4-m Blanco telescope at CTIO, we perform a search on 19,930 stars in the outer regions of the cluster. Though we find no clear planet detections, by combining our result with the upper limit enabled by R. L. Gilliland et al.’s (2000) Hubble search for planets around an independent sample of 34,091 stars in the inner cluster, we place the strongest limit to date on hot Jupiters with periods of $0.8\leq P\leq 8.3$ days and $0.5~R_{\rm Jup}\leq R_{\rm P}\leq 2.0~R_{\rm Jup}$ of $f_{\rm HJ}<0.11\%$, a factor of $\sim$}4 below the occurrence rate in the Kepler field. Our search found 35 transiting planet candidates, though we are ultimately able to rule out each without follow-up observations. We also found 4 eclipsing binaries, including 3 previously-uncataloged detached eclipsing binary stars.
journal: AJfacilities: Blanco (DECam)software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), CFITSIO (W. Pence, 1999), DoPHOT (M. Mateo & P. L. Schechter, 1989; P. L. Schechter et al., 1993), FITSH (A. Pál, 2012), ISIS (C. Alard & R. H. Lupton, 1998; C. Alard, 2000), JKTEBOP (J. Southworth, 2013), JKTLD (J. Southworth, 2008), matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), pandas (T. pandas development team, 2020), Source Extractor (E. Bertin & S. Arnouts, 1996), VARTOOLS (J. D. Hartman & G. Á. Bakos, 2016), WCSTOOLS (D. J. Mink, 1998; J. Mink, 2019)
1 Introduction
Though giant planet formation rates have been found to correlate with [Fe/H] (e.g., G. Gonzalez, 1997; N. C. Santos et al., 2004; D. A. Fischer & J. Valenti, 2005; J. A. Johnson et al., 2010), the specific elements responsible for the correlation have largely been unresolved. Volatiles and silicates are expected to be crucial ingredients for giant planet formation, because the silicate dust grains in protoplanetary disks provide surfaces for the more abundant volatiles to stick to, forming planetesimals (K. M. Pontoppidan et al., 2014; K. I. Öberg & E. A. Bergin, 2021; J. Nielsen et al., 2023). Beyond the ice line, these volatiles likely make up the majority of the protoplanetary disk’s solid mass. These compounds largely consist of $\alpha$-elements — elements formed in $\alpha$-capture reactions — such as oxygen, silicon, and magnesium, suggesting that the $\alpha$-abundance, $\alpha$/H] = [Fe/H] + $\alpha$/Fe], may influence planet occurrence rates more than [Fe/H]. However, our current sample of giant planet hosts prevents us from disentangling a relationship. Many come from Kepler and TESS, whose targets primarily lie within the high-[Fe/H], low-$\alpha$/Fe] thin disk (T. Bensby et al., 2003, 2005; B. E. Reddy et al., 2006; X. Guo et al., 2017; D. H. Weinberg et al., 2019; E. Griffith et al., 2021). Past observational limits for ground-based radial velocity and transit searches have similarly limited our sample to primarily disk stars. Some analyses of low-[Fe/H], high-$\alpha$/Fe] subsamples of stars in the Galactic thick disk such as V. Z. Adibekyan et al. (2012a, b) have found higher numbers of giant planets than are predicted by a [Fe/H] dependence, which could potentially be explained by the relative $\alpha$-enhancement. However, the bias of our giant planet sample towards disk stars prevents us from determining a conclusive explanation for this surplus: we lack observations in the full $\alpha$/Fe]-[Fe/H] parameter space, particularly in high-[Fe/H], high-$\alpha$/Fe] populations (e.g., R. F. Wilson et al., 2022).
Additionally, globular clusters allow us to search for planets in dense stellar environments — a region of parameter space that is largely unexplored. This parameter space is complicated by some theoretical uncertainty: we don’t yet know if, e.g., the gravitational interactions produced by these dense fields would inhibit or enhance hot Jupiter formation. A. S. Hamers & S. Tremaine (2017) posit that only the densest core regions of the clusters would eject all planets, and A. J. Winter et al. (2020) suggest that the gravitational perturbations in dense fields may initiate planet-planet scattering that actually enhances hot Jupiter formation. Other studies suggest that conditions in globular clusters beyond gravitational effects may adversely affect the likelihood of planet formation. T. A. Thompson (2013) theorizes that because dense regions such as globular clusters tend to form in hot starbursts, the high temperatures may evaporate the protoplanetary disks more quickly than giant planets can form. Further, E. J. Lee & P. F. Hopkins (2020) find that the ambient radiation in star forming regions is extremely high, which could potentially eliminate ice lines in protoplanetary disks entirely. Even if the ice line is not entirely eliminated, analyses such as A. J. Winter et al. (2022) find that in disks subjected to high far ultraviolet radiation, the growth and migration of giant planets is still suppressed. Thus, tighter observational constraints on planet occurrence rates in extreme environments are needed to constrain these seemingly contradictory theories.
Globular clusters are also of potential interest for planet searches because of their well-constrained, relatively advanced ages. For hot Jupiters in particular, studies such as D.-C. Chen et al. (2023) suggest that, even if hot Jupiters formed in globular clusters, the systems may be old enough that the planets have already been engulfed by their host stars due to orbital tidal decay. Improved constraints on hot Jupiter occurrence rates in globular clusters can thus improve our understanding of formation in this parameter space.
To date, planet searches of globular clusters have resulted in one successful detection in M4: a roughly 1.18 $R_{\text{Jup}}$ planet orbiting the millisecond pulsar PSR B1620-26 (H. B. Richer et al., 2003). Its location towards the core of M4 implies that planet formation in clusters may still be possible, despite their complex stellar dynamics (J. M. Fregeau et al., 2006). Other studies suggest it may indicate that — though globular clusters may have too low a metallicity for typical giant planet formation — they may be suitable for forming planets from supernova fall-back (e.g., M. E. Beer et al., 2004; S. Sigurdsson et al., 2008). However, no searches for globular cluster planets orbiting normal main sequence or giant stars have successfully discovered a planet, including a further search of M4 using K2 data (J. J. Wallace et al., 2020), which was limited to only $\sim$4,000 distinguishable stars due to Kepler’s 4”/px resolution.
@@ 1 Introduction
Surveys of 47 Tuc began with an HST search by R. L. Gilliland et al. (2000) (hereafter G00). Using the accepted hot Jupiter occurrence rate in the solar neighborhood for the time — 0.8-1.0% — they expected to find $\sim$17 planets in their observations. However, despite taking 8.3 days of near-continuous data for $\sim$34,000 stars, they observed no statistically significant transit events. In a later re-analysis using a simulated stellar sample based on giant planet hosting stars in the Kepler sample, K. Masuda & J. N. Winn (2017) (hereafter MW17) find that, if the occurrence rate of hot Jupiters in 47 Tuc was to match that of Kepler field stars, accounting for the host-mass dependence of giant planet occurrence rate (J. A. Johnson et al., 2010), G00 could only have expected to find $\sim$2 planets in their sample of 34,000 stars, with a $\sim$15% chance of finding no planets. Thus, the null result of the G00 survey is less statistically significant than initially thought.
@@
D. T. F. Weldrake et al. (2005) (hereafter W05) followed-up the G00 survey with a 33-night ground-based search using the Australian National University 40 inch Telescope. To complement the G00 search, which focused more on the core of the cluster, this survey focused on $\sim$22,000 stars in the outskirts of the cluster. Using Monte Carlo simulations, they expected to find $\sim$7 planets in their data if the occurrence rate was 0.8%, and still found none.
@@ 2.1 Survey Design
[…]
MISHAPS uses the Dark Energy Camera (DECam) on the 4-m Víctor M. Blanco telescope at Cerro-Tololo Interamerican Observatory to perform a high-cadence, wide-field survey of 47 Tuc and other $\alpha$-enhanced fields to search for transiting planets. DECam’s wide field of view and the Blanco’s 4-m diameter allow us to pursue an alternative strategy to most transiting planet searches that have relied in part on covering multiple transits and folding to increase the signal-to-noise ratio (S/N) above a detection threshold. Instead, the Blanco’s aperture enables a clear detection of a Jupiter-radius planet’s transit over $\sim$5 magnitudes of the 47 Tuc main sequence. We therefore planned a small survey ($\sim$7 nights) to search for single full or partial transits on a larger sample of stars than either G00 or W05. Without the necessity of multiple transit detections, the detection probability for short-period planets is therefore $\sim$10’s of percent for even 7 nights of data. With the expectation that the number of planet candidates we would find would be small, and that the quality of data would allow many EB false positives to be rejected from survey data alone, we anticipated that the number of targets that required follow-up would be small. For those that did, an estimate of the orbital period from a single transit would limit the amount of observations needed to recover a second transit and measure a period (e.g., S. Seager & G. Mallén-Ornelas, 2003; J. C. Yee & B. S. Gaudi, 2008; J. Villanueva et al., 2019).
We selected DECam for its size, photometric precision, near-infrared $z$-band sensitivity, and low overhead. DECam consists of 62 deep-depletion science CCDs, each of which has a size of 2048x4096 px for a total of 520 Mpx (K. Honscheid & D. L. DePoy, 2008), though only 60 are currently in use. The $\sim$3 deg2 FOV allows us to simultaneously monitor the full expanse of 47 Tuc. Its photometric precision allows us to resolve stars with magnitudes down to $r\approx 24$ and to potentially detect single transits on stars down to $r\approx 22$. DECam’s extremely high $z$-band sensitivity and low overheads are critical to enabling chromatic timeseries observations, which allow us to use differences in eclipse depths in $r$ and $z$ to rule out false positives caused by secondary eclipses of eclipsing binaries (EBs), diluted primary eclipses of hierarchical triple stars, and blended Small Magellanic Cloud (SMC) EBs. The low overheads also allow observations at a high enough cadence to resolve ingresses and egresses of transits, which are important for estimating accurate transit times and host stellar densities (e.g., S. Seager & G. Mallén-Ornelas, 2003), as well as increasing the accuracy with which one can estimate periods by adding well-sampled partial eclipses.
@@ 2.2 Observations
Preliminary observations of 47 Tuc were taken in short 2-4 hour windows on 17 nights in 2019 and 2021. Additionally, eight half-nights of observations of the cluster were conducted in 2022. Each night, we took a continuous timeseries of images with 100 s exposure times, alternating between the $r$ and $z$ filters with each exposure. This exposure sequence resulted in a $\sim 4$-minute cadence in each filter, or $\sim 2$ minutes combined. These sequences were interrupted roughly once per hour to check and correct the telescope pointing if it drifted by more than 10”. All images were first processed through the DECam Community Pipeline (F. Valdes et al., 2014), which applies calibration frames, corrects instrumental signatures, and applies a cosmic ray removal algorithm. We performed a preliminary determination of the overall quality of the nights by plotting the background, ellipticity, zeropoint magnitude, and seeing values from the image headers and flagging any nights which consistently had a zeropoint <<30.0, the magnitude of a star resulting in only 1 photon s-1 in the exposure. These poor-quality nights were later confirmed via analysis of the nightly median absolute deviation (MAD) of the lightcurve’s photometric root-mean-square (RMS) after the images were processed and the photometry was extracted (§4.1). Table 2.2 shows the log of observing nights and their mean data quality statistics. In total, our observations produce 1,288 images in the $z$-band and 1,293 images in the $r$-band.
@@ 3 Reference Image Analysis
We produce reference images for each band for use in difference imaging, requiring us to choose an appropriate number of high-quality images from our full dataset and stack them (§3.1). We then extract our initial catalog of stars using PSF photometry, and find astrometric solutions and photometric calibrations for each target by crossmatching with Gaia DR3 (C. Babusiaux et al., 2023) and the NOIRLab Source Catalog Data Release 2 (NSC, D. L. Nidever et al., 2018, 2021)3.2). We also determine appropriate proper motion and color selections for cluster membership using the Gaia crossmatch results (§3.3). We then perform several characterization steps to determine which targets will be used to perform transit injections and for use in efficiency estimates (§3.4).
@@ 3.1 Reference Image Construction
For each of DECam’s CCD chips, we constructed a deep, high-quality reference image in the $r$ and $z$ filters by selecting and combining a small number of the best images from the data taken in 2021, which had the best conditions. To do this, the community pipeline-determined point spread function (PSF), full width at half maximum (FWHM), sky background, zeropoint (magnitude resulting in 1 photon s-1), PSF ellipticity, and airmass for all images were compiled from the FITS headers, and sorted by FWHM. Fifteen images were selected in order of increasing FWHM, and any with high sky background, PSF ellipticity, or smaller zeropoint magnitude were skipped. Each candidate image was visually inspected to ensure it was a good-quality image. The fifteen selected images were combined using the ISIS image subtraction program’s ref.csh task (C. Alard & R. H. Lupton, 1998; C. Alard, 2000), which aligns the images; determines a PSF-matching convolution kernel and photometric scaling to match the PSF, zeropoint, and sky background of the input images; and then combines them as a mean stack with 3-$\sigma$ clipping. We refer to the resultant images as the reference image for the relevant chip and filter.